Vin Diesel owns the Blue Spruce Barber Shop. He employs 5 barbers and pays each a base rate of $1,490 per month. One of the barbers serves as the manager and receives an extra $550 per month. In addition to the base rate, each barber also receives a commission of $4.50 per haircut.

Other costs are as follows.
Advertising $270 per month
Rent $990 per month
Barber supplies $0.30 per haircut
Utilities $185 per month plus $0.20 per haircut
Magazines $20 per month

Vin currently charges $10.00 per haircut.

Collapse question part

(a)

Correct answer. Your answer is correct.
Determine the variable costs per haircut and the total monthly fixed costs. (Round variable costs to 2 decimal places, e.g. 2.25.)
Total variable cost per haircut $

Entry field with correct answer

Total fixed $

Entry field with correct answer

Click if you would like to Show Work for this question:

Open Show Work

SHOW SOLUTION

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

Attempts: 1 of 3 used

Collapse question part

(b)

Compute the break-even point in units and dollars. (Round answers to 0 decimal places, e.g. 1,225.)
Break-even point in units

haircuts
Break-even point in dollars $

Click if you would like to Show Work for this question:
Answer #1

Variable cost = 4.50+0.30+0.20 = 5

Fixed cost = (5*1490)+550+270+990+185+20

= 8915

B)

Breakeven units = Fixed cost/Contribution margin per unit

= 8915/(10-5)

= 1783

Breakeven Dollars = 1783*10

= 17,830